y+4=-4(x-2)
rewrite this equation in slope-intercept form?
rewrite this equation in standard form?
what is the equation in standard form of perpendicular line that passes through(5,1)
what is the x-intercept of the perpendicular line
thank you

Answer by rothauserc(4718)

y +4 = -4(x-2)
****************************************************************************
slope intercept form
y = -4x +8 -4
y = -4x +4
***************************************************************************
standard form
4x +y = 4
***************************************************************************
slope of line is -4
slope of perpendicular line is 1/4 (this is the negative reciprocal of the line's slope)
y = x/4 + b
now use the point (5, 1)
1 = 5/4 + b
b = -1/4
y = x/4 - 1/4
-x/4 +y = -1/4
multiply both sides of = by -4
standard form of perpendicular line is
x -4y = 1
***************************************************************************
x = 1 +4y
if y=0, then x=1
the x intercept is at x=1 and the point of intersection is (1, 0)